OpenStudy (anonymous):

The peremeter of a rectangle pool is 140 yd. If the L is 10 yd more than the W , find the L + W.

OpenStudy (anonymous):

L+W=70

OpenStudy (anonymous):

easy question see here given that L = W+10yd perimeter of rectangle = 2(L+B) hence here 2(L+B) = 2(W+10yd. + W) 2(2W+10yd)=140yd divide both sides by2 2W+10yd=140yd/2 2W+10yd = 70yd 2W = 70yd - 10yd 2W = 60 yd. hence W = 60/2 yd. W = 30 yards now L = 10 + W L = 40yds. ( yards) hence W + L = 70yards it can be solved much easily but i solved it like this so that u can understand how to values of L and B also got it or not